OpenStudy (josee):

How close to −3 do we have to take x for the inequality 1/(x + 3)^2 > 700 to hold? 1. within at least 0.097 2. within at least 0.057 3. within at least 0.037 4. within at least 0.077 5. within at least 0.017

OpenStudy (anonymous):

x has to be within \[\sqrt{\frac{1}{700}}\] of -3 ,whatever that number is

OpenStudy (anonymous):

.037 is the nearest

OpenStudy (anonymous):

need the method?

OpenStudy (josee):

oh so the -3 has nothing t do with the problem? it's basically seeing if it falls on that range?

OpenStudy (anonymous):

oh it has something all right flip the inequality to get \[(x+3)^2<\frac{1}{700}\] take the square root get \[x+3<\frac{1}{\sqrt{700}}\]

OpenStudy (anonymous):

well that is a mistake sorry lets do it right

OpenStudy (anonymous):

\[(x+3)^2<\frac{1}{700}\] \[-\frac{1}{\sqrt{700}}<x+3<\frac{1}{\sqrt{700}}\]

OpenStudy (anonymous):

or \[|x+3|<\frac{1}{\sqrt{700}}\]

OpenStudy (anonymous):

which is the math way to say x is within .037 of -3

OpenStudy (josee):

gotcha!

OpenStudy (anonymous):

go ahead and post i will take a look

OpenStudy (josee):

im having trouble with 5,6,7,11,12,13,14 :/

OpenStudy (anonymous):

that is a lot of troubles! i will do 5 first. the function \[g(x)=\frac{2x-5}{|x-7|}\] has a vertical asymptote at x = 7. i am not sure why it says it is a "jump" discontinuity, but never mind.

OpenStudy (anonymous):

as you approach from the right it goes to positive infinity. f(7) = 0 but it is positive on the right. so that limit will be positive infinity

OpenStudy (anonymous):

\[f(x)=\frac{x^2-x-20}{x^2-7x+10}\]

OpenStudy (josee):

its not positive infinity :/

OpenStudy (anonymous):

hold on let me reread the problem, but let me finish #6

OpenStudy (anonymous):

\[f(x)=\frac{(x-5)(x+4)}{(x-5)(x-2)}=\frac{x+4}{x-2}\]

OpenStudy (anonymous):

you can't get rid of the discontinuity at 2

OpenStudy (anonymous):

if you replace x by 5 you get \[\frac{9}{3}\] which is not \[\frac{10}{3}\] so the function is not continuous at 5 either

OpenStudy (anonymous):

#7 C holds for sure, because the function is continuous on [-2,2] and negative at -2 , positive at 2, so it must cross the x axis somewhere B is silly, you have no idea what the function looks like. \[f(-2)+(-2)=-1, f(2)+2=1\] so \[f(x)+x\] must also cross the x axis

OpenStudy (anonymous):

so A and C

OpenStudy (anonymous):

if it is not positive infinity what is it?

OpenStudy (josee):

idk:/ im thinking it doesnt exist

OpenStudy (anonymous):

hmmm i don't know why. hold on

OpenStudy (anonymous):

that is the one with the graph right?

OpenStudy (anonymous):

oh crap i was wrong. sorry. does not exist is right i believe my mistake

OpenStudy (anonymous):

#11 at 1:00 accelerating because slope of tangent lines are increasing. as for first part in 10 minutes he went 40 km

OpenStudy (anonymous):

so that is exactly 120 km/hr average, not less these are tricky!

OpenStudy (anonymous):

where did you get these?

OpenStudy (josee):

Lol my teacher sends them to me for a review.

OpenStudy (anonymous):

12 is actually easiest of all

OpenStudy (anonymous):

is your teacher uri treismann?

OpenStudy (josee):

yess

OpenStudy (anonymous):

lol

OpenStudy (josee):

omg number 5 isnt that the limit doesnt exist lol

OpenStudy (anonymous):

wtf

OpenStudy (anonymous):

what is it???

OpenStudy (josee):

omg it isnt -7..

OpenStudy (josee):

i have to submit my answer and it tells me if its wrong or not. if it is, i gotta try again

OpenStudy (anonymous):

let me post this damned problem and see if we get a response. but first lets do 12

OpenStudy (anonymous):

ready?

OpenStudy (anonymous):

the derivative is the slope of the tangent line. if you look at the picture you will see that at -2 the function is going down, so \[g'(2)<0\]

OpenStudy (josee):

it was 9. and 11 was B only

OpenStudy (anonymous):

at 0, 1 and 3 it is increasing , so we know that the order begins \[g'(2),0,...\]

OpenStudy (anonymous):

yes we knew 11 was B only. maybe my answer was not clear. it was exactly 120 km/hr , not less than

OpenStudy (anonymous):

9???!!!

OpenStudy (anonymous):

hold on let me post it. but 12 should be easy to finish. the derivative is the slope of the tangent line, so arrange so that slopes go from smallest to biggest

OpenStudy (anonymous):

#6 is correct for problem 12

OpenStudy (josee):

i just need number 13 :D